OpenStudy (study312):

If a sprinkler waters 1 over 12 of a lawn in 1 over 2 hour, how much time will it take to water the entire lawn?

OpenStudy (study312):

A. 7/12 B. 1/6 C. 6 D. 10

OpenStudy (study312):

I think it's 6 because flip 1/2 to 2/1 the 1 • 2 = 2 and 12 • 1 = 12 which would equal 2/12. I see that 2 and 12 remind me of 6 so 12 ÷ 2 = 6

OpenStudy (study312):

Am I right?

OpenStudy (phi):

There are a few ways to do this. one way is write a ratio \[ \frac{ \frac{1}{12} \ lawn}{\frac{1}{2} \ hour}= \frac{1 \ lawn}{x \ hours} \]

OpenStudy (phi):

the other way to think of it is as rate * time = 1 lawn rate is the amount of lawn watered divided by time time is unknown, call it x so you get \[ \frac{ \frac{1}{12} \ lawn}{\frac{1}{2} \ hour} x=1 \]

OpenStudy (phi):

in either case, you should simplify the messy fraction \[ \frac{ \frac{1}{12} \ lawn}{\frac{1}{2} \ hour} \] by mutliplying top and bottom by 2 \[ \frac{ \frac{1}{12} \cdot 2}{\frac{1}{2} \cdot 2} \] 1/2 * 2 is 1 so the fraction becomes \[ \frac{ \frac{1}{12} \cdot 2}{1}= \frac{1}{12} \cdot 2 \] and that becomes \[ \frac{2}{12} = \frac{1}{6} \]

OpenStudy (phi):

now you can solve for the time 1/6 * x = 1 multiply both sides by 6 you get x=6 hours or using ratios \[ \frac{1}{6}= \frac{1}{x}\] cross multiply to get x=6